OpenStudy (anonymous):

Compute |v1| and |v2| where v1=(2,-6) and v2=(-4,7)

OpenStudy (anonymous):

I assume that these are vectors?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

if z is the vector z = <a,b>, then |z| = sqrt(a^2 + b^2)

OpenStudy (anonymous):

Remember that if \(\mathbf{v} = \langle a,b\rangle\), then \(\|\mathbf{v}\| = \sqrt{a^2+b^2}\).

OpenStudy (anonymous):

So |v1| would be the square root of 40?

jimthompson5910 (jim_thompson5910):

Correct. But I would simplify that.

OpenStudy (anonymous):

Right. My thanks.

jimthompson5910 (jim_thompson5910):

What do you get when you simplify \(\Large \sqrt{40}\)